# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ VI Conclusion & Future Works
In this work, we address a critical challenge in object detection: risk control. While the topic had previously been studied in some restricted cases (for a single model, or a sub-task of OD), we provide the first framework to build guaranteed prediction sets for all common tasks of object detection.
@@
Our main contribution lies in the introduction of Sequential Conformal Risk Control (SeqCRC). Unlike previous works using CRC, SeqCRC allows to control risks of multiple dependent parameters, enabling simultaneous control of confidence thresholds, classification score and localization margin.
We then study different ways to measure prediction errors (conformal loss functions) – reviewing existing ones and proposing new ones suitable for object detection. We discuss their properties, and also consider how best to match predicted objects with the real ones (ground truths).
We test our framework through experiments, confirming that our SeqCRC method functions correctly and efficiently. A key outcome is the creation of the first benchmark for Conformal Object Detection, establishing a standardized way to compare methods in this field. Our analysis also highlights the strengths and weaknesses of this approach, identifying areas that require further research.
@@
To support the research community, we provide an open-source codebase used to generate our benchmark. This resource includes models and procedures, enabling development of new techniques, error metrics, or guarantees based on this work.
@@
Our analysis identifies key directions for future research. One direction is exploring less conservative approaches for obtaining global guarantees than the current one, which could lead to tighter prediction sets. Another area is the development of loss functions that better balance valuable guarantees with smaller, i.e. more informative, prediction set sizes. Extending the framework to control false positives (precision) and improving its robustness against inconsistencies in ground-truth annotations present additional challenges. Further optimization of component choices, guided by our benchmark and supported by our open-source codebase, will enhance practical applicability.
Overall, this research provides a solid foundation for building increasingly trustworthy object detection systems. We offer a novel method (SeqCRC) and valuable resources (the benchmark and codebase) to advance the development of reliable computer vision applications.
Appendix A Level of Guarantee, Aggregation & Matching
Level of Guarantee and Aggregation Functions:
On a given image, there can a priori be any number of objects (and their associated bounding boxes). On the one hand, each of those individual objects may be considered as ground truth individually. On the other hand, the ground truth can be defined as the set of all objects present on an image. Formally, the first setting corresponds to $Y_{i}=(b_{i},c_{i})$, while the latter corresponds to the setting introduced in the article, i.e. $Y_{i}=(b_{i,j},c_{i,j})_{j=1}^{n_{i}}$. Respectively, we refer to those as object-level and image-level guarantees following [14, 16, 15]. Object-level guarantees have previously been referred to as box-level in [14, 16]. However, the exact setting varied, as conformal prediction was only applied to true-positive bounding boxes (i.e., ground-truth - prediction pairs, with only pairs with IoU above a certain threshold $\beta$ considered).
The choice depends on practical requirements and should be reflected in the definition of the losses. An example is a binary loss: in the object-level setting, it entails that a proportion at least $1-\alpha$ of ground truth bounding boxes are correctly covered, while in the image-level setting, it entails that, on a proportion $1-\alpha$ of images at least, all bounding boxes (or a desired proportion) are covered.
Matching:
As common in object detection [40, 43], a matching between the true bounding boxes and predicted ones is necessary, not only for evaluation here but also for calibration. We refer to by matching, a family of functions $\pi_{i}$ defined for each image $X_{i}$ that maps the index $j$ of an object on $X_{i}$ to the index of a predicted box among $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$. More formally, $\pi_{i}:\{1,\cdots,n_{i}\}\rightarrow\{1,\cdots,\lvert\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})\rvert\}$ with $n_{i}$ the number of objects on image $X_{i}$ and $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(X_{i})$ the sets of bounding boxes likely to correspond to objects on image $X_{i}$. We do not make further assumption on the matching algorithm. Therefore it includes injective functions, i.e. each prediction being matched to at most one ground truth. We can also consider functions where some inputs might have identic output, i.e. reusing predictions for multiple ground truth. An example of an injective matching is the commonly used Hungarian algorithm. Notably, allowing reusing a same predicted bound for more than one ground truth object is an approach to avoid the true-positive limit of [14].
Appendix B Proofs
@@ B-A Main theorem
Suppose that Assumptions 1 and 3 hold true. [^18] See Footnote 10. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
@@
+=$\mathcal{L}_{+}^{\bullet}=\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{nR^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})}{n+1}+\frac{B^{\bullet}}{n+1}\leq\alpha^{\bullet}\right\}$ (18)
@@
On the other hand, if $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, then by definition of $\lambda^{\mathrm{cnf}}_{-}$, [^19] In particular, when $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, the set in (5) is nonempty. by $R^{\bullet}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\bullet})\leq\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$, and by right-continuity of $R^{\bullet}_{n}$ in its first argument, we have $\frac{n}{n+1}R^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\bar{\lambda}^{\bullet})\leq\alpha^{\text{cnf}}$ almost surely. Combining this with the assumption $\alpha^{\bullet}\geq\alpha^{\text{cnf}}+\frac{B^{\bullet}}{n+1}$ implies that $\bar{\lambda}^{\bullet}\in\mathcal{L}_{+}^{\bullet}$ as in the previous case.
@@
First note that $\tilde{R}^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely. (Indeed, by assumption, $R^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ and $R^{\bullet}_{n+1}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ a.s.) Therefore, the set in (19) is nonempty and $\lambda^{\mathrm{cnf}}_{*}$ is well defined. [^20] In Step 3.a below, we will also handle cases in which this set is empty. Moreover, by (19), by $R^{\mathrm{cnf}}_{n+1}\leq\tilde{R}^{\mathrm{cnf}}_{n+1}$ and by right-continuity of $R^{\mathrm{cnf}}_{n+1}$, we have $R^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}}_{*})\leq\alpha^{\text{cnf}}$ almost surely. Taking the expectation and noting that $\lambda^{\mathrm{cnf}}_{*}$ is symmetric (i.e., invariant under any permutation of the $n+1$ calibration points; see Lemma 1 in Appendix B-B below) yields $\mathbb{E}\left[L_{n+1}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{*})\right]\leq\alpha^{\mathrm{cnf}}$. The fact that $\lambda^{\mathrm{cnf}}_{+}\geq\lambda^{\mathrm{cnf}}_{*}$ (see details in Step 3.a below) and that $L_{n+1}^{\mathrm{cnf}}$ is non-increasing concludes the proof of (8).
Step 3: proof of Eq. (7). The rest of the proof is novel. Though similar in spirit to that of [53, Theorem 1], our multitask extension requires some additional care because $\lambda^{\bullet}_{+}$ uses the same calibration data as $\lambda^{\mathrm{cnf}}_{+}$. This is in contrast with, e.g., the analysis of [24, Remark 1 and Theorem 1] that works for two independent calibration datasets (data-splitting trick). Note moreover that, contrary to Step 2, in all the sequel, we do not assume that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i$.
@@
Step 3.a: we show that $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely. To that end, recall that [^21] Though not written explicitly to simplify notation, the parameter $\lambda^{\mathrm{cnf}}$ runs over $\Lambda^{\mathrm{cnf}}$ in each of the three sets below.
@@
Note that the three sets above, say, $A_{+}\subset A_{*}\subset A_{-}$, are nested. (This is by definition of $\tilde{R}^{\mathrm{cnf}}_{n+1}$ in (20) and because $L^{\mathrm{cnf}}_{n+1}$ and $L^{\bullet}_{n+1}$ are bounded between 00 and $\tilde{B}^{\mathrm{cnf}}=\max\{B^{\mathrm{cnf}},B^{\mathrm{loc}},B^{\mathrm{cls}}\}$ by Assumption 3.). Taking the infimum (with the convention $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$) yields $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely.
@@
Step 3.c: we show that $\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]\leq\alpha^{\bullet}$, by using the fact that $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ are symmetric in the sense of Lemma 1 in Appendix B-B below.
@@
B-B A simple symmetry property
[…]
@@ B-B A simple symmetry property
$\lambda^{\mathrm{cnf}}_{*}\left((L^{\text{cnf}}_{\sigma(1)},L^{\text{loc}}_{\sigma(1)},L^{\text{cls}}_{\sigma(1)}),\ldots,(L^{\text{cnf}}_{\sigma(n+1)},L^{\text{loc}}_{\sigma(n+1)},L^{\text{cls}}_{\sigma(n+1)})\right)$ (25)
@@
$\lambda^{\bullet}_{*}\left((L^{\text{cnf}}_{\sigma(1)},L^{\text{loc}}_{\sigma(1)},L^{\text{cls}}_{\sigma(1)}),\ldots,(L^{\text{cnf}}_{\sigma(n+1)},L^{\text{loc}}_{\sigma(n+1)},L^{\text{cls}}_{\sigma(n+1)})\right).$ (27)

### v2

@@ VI Conclusion & Future Works
In this work, we address the critical challenge of building trustworthy object detection (OD) systems. We focus on a family of statistically valid uncertainty quantification methods: conformal risk control. Though this topic had previously been studied in some restricted OD cases (for a single model, or a sub-task of OD), we provide the first framework to build guaranteed prediction sets for all common tasks in OD.
@@
Our main contribution is the new statistical method Sequential Conformal Risk Control (SeqCRC). Unlike previous works using CRC, SeqCRC allows to control risks associated with two sequential tasks using two dependent parameters. In OD this provides a statistically sound way to tune a confidence threshold as well as classification and localization “error margins”.
We study various ways to measure prediction errors (conformal loss functions), both old and new, and discuss their properties as well as other key hyperparameters (e.g., matching distance). We test our framework through extensive experiments, which confirm that our SeqCRC method effectively controls different risks and produces small enough set sizes.
@@
To support the research community, we provide an open-source codebase that we used for our experiments. It includes models and procedures, enabling development of new techniques, error metrics, and associated guarantees. We hope this codebase and our extensive experiments can be used for future benchmarks in Conformal Object Detection.
@@
Our analysis identifies key directions for future research. For instance, designing new confidence loss functions that better balance valuable guarantees with smaller (more informative) prediction sets would be key in practice. Other examples mentioned in Section V-D are about extending the framework to control false positives (precision), or improving its robustness against inconsistencies in ground-truth annotations.
Acknowledgements
Our work has benefitted from the AI Cluster ANITI and the research program DEEL. ANITI is funded by the France 2030 program under the Grant agreement n°ANR-23-IACL-0002. DEEL is an integrative program of the AI Cluster ANITI, designed and operated jointly with IRT Saint Exupéry, with the financial support from its industrial and academic partners and the France 2030 program under the Grant agreement n°ANR-10-AIRT-01.
A. Mazoyer has been supported by the project ROMEO (ANR-21-ASIA-0001) from the ASTRID joint program of the French National Research Agency (ANR) and the French Innovation and Defence Agency (AID)
Appendix A Proofs
@@ A-A Main theorem
Suppose that Assumptions 1 and 3 hold true.See Footnote 2. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
$\mathbb{E}\left[L_{n+1}^{\bullet}(\lambda^{\mathrm{cnf}}_{+},\lambda^{\bullet}_{+})\right]\leq\alpha^{\bullet}\;.$ (7)
@@
$\mathbb{E}\left[L_{n+1}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{+})\right]\leq\alpha^{\mathrm{cnf}}\;.$ (8)
@@
$\mathcal{L}_{+}^{\bullet}=\left\{\lambda^{\bullet}\in\Lambda^{\bullet}:\frac{nR^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\lambda^{\bullet})}{n+1}+\frac{B^{\bullet}}{n+1}\leq\alpha^{\bullet}\right\}$ (18)
@@
On the other hand, if $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, then by definition of $\lambda^{\mathrm{cnf}}_{-}$,In particular, when $\lambda^{\mathrm{cnf}}_{-}<\bar{\lambda}^{\mathrm{cnf}}$, the set in (5) is nonempty. by $R^{\bullet}_{n}(\lambda^{\mathrm{cnf}},\bar{\lambda}^{\bullet})\leq\tilde{R}^{\mathrm{cnf}}_{n}(\lambda^{\mathrm{cnf}})$, and by right-continuity of $R^{\bullet}_{n}$ in its first argument, we have $\frac{n}{n+1}R^{\bullet}_{n}(\lambda^{\mathrm{cnf}}_{-},\bar{\lambda}^{\bullet})\leq\alpha^{\text{cnf}}$ almost surely. Combining this with the assumption $\alpha^{\bullet}\geq\alpha^{\text{cnf}}+\frac{B^{\bullet}}{n+1}$ implies that $\bar{\lambda}^{\bullet}\in\mathcal{L}_{+}^{\bullet}$ as in the previous case.
@@
First note that $\tilde{R}^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely. (Indeed, by assumption, $R^{\mathrm{cnf}}_{n+1}(\bar{\lambda}^{\mathrm{cnf}})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ and $R^{\bullet}_{n+1}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=\frac{1}{n+1}\sum_{i=1}^{n+1}L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ a.s.) Therefore, the set in (19) is nonempty and $\lambda^{\mathrm{cnf}}_{*}$ is well defined.In Step 3.a below, we will also handle cases in which this set is empty. Moreover, by (19), by $R^{\mathrm{cnf}}_{n+1}\leq\tilde{R}^{\mathrm{cnf}}_{n+1}$ and by right-continuity of $R^{\mathrm{cnf}}_{n+1}$, we have $R^{\mathrm{cnf}}_{n+1}(\lambda^{\mathrm{cnf}}_{*})\leq\alpha^{\text{cnf}}$ almost surely. Taking the expectation and noting that $\lambda^{\mathrm{cnf}}_{*}$ is symmetric (i.e., invariant under any permutation of the $n+1$ calibration/test points; see Lemma 1 in Appendix A-B below) yields $\mathbb{E}\left[L_{n+1}^{\mathrm{cnf}}(\lambda^{\mathrm{cnf}}_{*})\right]\leq\alpha^{\mathrm{cnf}}$. The fact that $\lambda^{\mathrm{cnf}}_{+}\geq\lambda^{\mathrm{cnf}}_{*}$ (see details in Step 3.a below) and that $L_{n+1}^{\mathrm{cnf}}$ is non-increasing concludes the proof of (8).
Step 3: proof of Eq. (7). The rest of the proof is novel. Though similar in spirit to that of [53, Theorem 1], our multitask extension requires some additional care because $\lambda^{\bullet}_{+}$ uses the same calibration data as $\lambda^{\mathrm{cnf}}_{+}$. This is in contrast with, e.g., the analysis of [25, Theorem 3 in Section 3.2] that works for two independent calibration datasets (data-splitting trick). Note moreover that, contrary to Step 2, in all the sequel, we do not assume that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i$.
@@
Step 3.a: we show that $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely. To that end, recall thatThough not written explicitly to simplify notation, the parameter $\lambda^{\mathrm{cnf}}$ runs over $\Lambda^{\mathrm{cnf}}$ in each of the three sets below.
@@
Note that the three sets above, say, $A_{+}\subset A_{*}\subset A_{-}$, are nested. (This is by definition of $\tilde{R}^{\mathrm{cnf}}_{n+1}$ in (20) and because $L^{\mathrm{cnf}}_{n+1}$ and $L^{\bullet}_{n+1}$ are bounded between 0 and $\tilde{B}^{\mathrm{cnf}}=\max\{B^{\mathrm{cnf}},B^{\mathrm{loc}},B^{\mathrm{cls}}\}$ by Assumption 3.). Taking the infimum (with the convention $\inf\varnothing=\bar{\lambda}^{\mathrm{cnf}}$) yields $\lambda^{\mathrm{cnf}}_{-}\leq\lambda^{\mathrm{cnf}}_{*}\leq\lambda^{\mathrm{cnf}}_{+}$ almost surely.
@@
Step 3.c: we show that $\mathbb{E}\left[L^{\bullet}_{n+1}(\lambda^{\mathrm{cnf}}_{*},\lambda^{\bullet}_{*})\right]\leq\alpha^{\bullet}$, by using the fact that $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$ are symmetric in the sense of Lemma 1 in Appendix A-B below.
@@
A-B A simple symmetry property
We now state and prove a simple symmetry property that is key to obtaining Theorem 2. Put briefly, it means that the oracle parameters $\lambda^{\mathrm{cnf}}_{*}$ and $\lambda^{\bullet}_{*}$, which are functions of $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$, are invariant under any permutation of these loss $3$-tuples. Recall that we use the placeholder $\bullet\in\{\text{loc},\text{cls}\}$.
@@ A-B A simple symmetry property
[…]
@@
[…]